# Online Smoothed Demand Management

### v1

@@ 6.1 The PALD Framework
To tackle the above challenges, rather than defining a new prediction model and designing an algorithm around it, we propose a new learning framework (Partitioned Accounting & Learned Decisions (PALD)) that directly “learns the best algorithm” within a general class that guarantees a certain worst-case robustness.
@@
Motivation. We take insight from the competitive proof of PAAD (Theorem 4.1): under the pseudo-cost minimization framework, a key step in bounding the competitive ratio uses characteristics of the threshold functions ($\phi_{b}$ for base drivers, $\phi_{f}$ and $\psi_{f}$ for flexible drivers) that govern the algorithm’s decisions. In particular, the definition of the pseudo-cost minimization problem allows us to relate PAAD’s cost to that of (a lower bound on) the offline optimal cost via these threshold functions. Thus, for a given target robustness $\rho>\alpha$ (where $\alpha$ is the competitive ratio of PAAD for OSDM-S), we can exactly define a feasible set of threshold functions, based on the characteristics used in the competitiveness proof, that guarantee $\rho$-robustness for PALD. A similar characterization of a “robust threshold set” was derived by Benomar:25 for the simpler one-max search problem to investigate the design of algorithms’ “smoothness”; however, in PALD, our goal is instead to directly learn the best threshold functions (denoted by $\hat{\phi}_{b},\hat{\phi}_{f}$ and $\hat{\psi}_{f}$) within this feasible set to optimize a downstream metric (e.g., empirical competitive ratio) on historical instances. Given the learned thresholds $\hat{\phi}_{b},\hat{\phi}_{f}$ and $\hat{\psi}_{f}$, PALD’s pseudocode is exactly the same as that of PAAD (see Algorithm 2) but with $\hat{\phi}_{b},\hat{\phi}_{f}$ and $\hat{\psi}_{f}$ in place of the (analytical) thresholds $\phi_{b},\phi_{f}$ and $\psi_{f}$.
@@
$\Phi_{b}(0$ $,w)+(1-w)(p_{\max}+2\gamma)+p_{\max}(c+\varepsilon)+2\delta-cwp_{\min}\leq\rho\left[\phi_{b}(w)-2\gamma+\varepsilon p_{\max}+\nicefrac{{2\kappa}}{{T}}\right]\big\}.$
@@
$\hskip 41.62473pt\Psi_{f}(0,v)+(1\!-\!v)(p_{\max}(c\!+\!\varepsilon)\!+\!2\delta)\leq\rho\big[\nicefrac{{1}}{{\omega}}\big(\phi_{f}(w)+\psi_{f}(v)-2\kappa\big)+\nicefrac{{2\kappa}}{{T}}\big]\big\}.$
@@
Proof Sketch of Theorem 6.1. (full proof in § G.1) Since PALD is identical to PAAD save for the learned threshold functions, we inherit feasibility from Lemma E.1. Then, we give lemmas to bound the costs of OPT and PALD, and the relationship between the two. First, we lower bound the cost of OPT; as in the proof of Theorem 4.1, to do this we partition the time horizon into active and inactive periods. In the case of PALD, since the partitioning relies on the value of the base driver threshold function and PALD’s is learned, we must also use the definition of $\mathcal{R}_{b}$ to prove that OPT’s cost during inactive periods can be bounded in the same manner. Next, we upper bound the cost of PALD by “charging” each driver’s decisions to their corresponding threshold functions and accounting for any worst-case purchasing, switching, and delivery costs that arise due to constraints for a particular unit of demand. We show that the same charging argument from Theorem 4.1 still holds with learned thresholds. Finally, we use the definitions of $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ to relate the cost of PALD to that of OPT. Using a similar contradiction argument as in the proof of Theorem 4.1, the last inequalities in the definition of the feasible sets are exactly those needed to show that the ratio $\frac{\texttt{PALD}(\mathcal{I})-p_{\max}\hat{s}}{\texttt{OPT}(\mathcal{I})}$ must be at most $\rho$, completing the proof. $\square$
@@
Remark. PALD exhibits some similarities with existing approaches for data-driven algorithms: for instance, Zeynali:21:DataDriven learn the best exponent for an exponential threshold function in the online knapsack problem—our approach is distinguished by its expressiveness in allowing any arbitrary threshold function while retaining worst-case guarantees. Li:22:ExpertCalibrated learn an ML model to solve online convex optimization that is calibrated (although not guaranteed) to be robust with respect to a baseline—in contrast, PALD allows us to specify the worst-case level performance we are willing to incur, and to learn an algorithm that performs well subject to that constraint with certainty.
6.2 Learning Methodology
@@
Using the PALD framework outlined above, we describe the learning methodology we use in our case study (see § 7) to learn robust threshold functions that minimize the empirical competitive ratio on historical instances while guaranteeing robustness. We parameterize each threshold function as a piecewise-affine function using a “knot” representation at a fixed grid of $K>1$ points $\{0,\nicefrac{{1}}{{K-1}},\nicefrac{{2}}{{K-1}},\ldots,1\}$. We learn a parameter vector $\mathbf{y}\in\mathbb{R}_{+}^{K}$ where each element defines the function’s value at a knot (between knots, the function is interpolated linearly). We denote by $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ the parameter vectors for $\smash{\hat{\phi}_{b},\hat{\phi}_{f},\hat{\psi}_{f}}$, respectively. This makes the robust sets tractable:
@@
The full proof is in § G.3. The key idea is that the feasibility conditions can be expressed as box constraints, linear inequalities, and affine halfspaces, whose intersection is a convex set.
@@
Recall that the analytical thresholds $\phi_{b},\phi_{f}$ and $\psi_{f}$ (Defs. 4.1 & 4.2) are continuous and monotone non-increasing functions—by definition of the PAAD algorithm, these lie in the feasible sets $\mathcal{R}_{b}(\alpha)$ and $\mathcal{R}_{f}(\alpha)$, respectively. For $\rho$ increasing away from $\alpha$, the feasible sets $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ become strictly larger, since the robustness conditions become easier to satisfy. Since piecewise-affine functions are known to be dense in continuous functions Cheney:1998:Density; Pinkus:2005:Density, the analytical thresholds can be approximated arbitrarily well using piecewise-affine functions with sufficiently many knots. Thus, for a given robustness $\rho>\alpha$, the feasible sets $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ are non-empty since they at least contain the analytical thresholds. Lemma 6.3 implies that it is efficient to project any learned piecewise-affine function into the feasible sets—this makes robustness tractable to enforce during learning, and unlocks several learning techniques which we discuss below.
@@
Directly Learning Thresholds. We implement the PALD algorithm by using differentiable convex optimization layers (e.g., CVXPYLayers CVXPYlayers:19) to solve the pseudo-cost minimization problem at each time step. This allows backpropagation through the entire PALD procedure to minimize the empirical competitive ratio on historical instances. Given a data set of historical instances $\mathcal{D}=\{\mathcal{I}_{1},\mathcal{I}_{2},\ldots,\mathcal{I}_{N}\}$, we can use this differentiable implementation of PALD to find a single set of parameters $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ that minimizes the empirical competitive ratio on $\mathcal{D}$, using projected gradient descent (PGD) to ensure that, during training, the learned thresholds lie in the robust sets $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ for a target $\rho>\alpha$. We refer to this as PALD-S (“S” for simple) in our case study (§ 7).
@@
End-to-End Contextual Learning. While directly learning a single set of thresholds improves average-case performance, it cannot leverage instance-specific context that is available and extremely useful in practice. For instance, our case study in § 7 uses locational marginal price (LMP) data for different grids; LMPs are known to correlate with factors such as the time of day, the weather, and the season Qu:24. Using a learning method that is aware of such context can enable better performance in practice by learning thresholds that are tailored to likely operating conditions.
To realize this context-aware approach, we employ a neural network $F_{\Theta}(\mathbf{x})\to\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ (with parameters $\Theta$) that takes as input a context vector $\mathbf{x}$ about a given driver and outputs predicted parameter vectors $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ that define the threshold functions. The final layer of this neural network is a differentiable projection layer that projects the output parameters $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ into the feasible sets $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ to ensure robustness at a target level $\rho>\alpha$.
@@
We construct an unsupervised learning pipeline that simulates the differentiable PALD algorithm on a historical data set $\mathcal{D}$. For each instance, each new driver in PALD makes a forward pass through the neural network $F_{\Theta}$ to obtain its context-aware thresholds $\smash{\hat{\phi}_{b},\hat{\phi}_{f}}$, and $\smash{\hat{\psi}_{f}}$. After observing the resulting performance of PALD on the data set, we backpropagate through the entire pipeline to learn a better neural network mapping between context and thresholds. We refer to this technique as PALD-C in our case study (§ 7), and give a diagram of it in Appendix A.3.
@@
7 Case Study: A Grid-integrated Data Center with Local Energy Storage
@@
Table 1: Statistics for LMP traces. All traces span the time period 01/01/2024-12/31/2024 at a 15-minute granularity, and LMP data is reported in \$ (USD) per megawatt-hour (MWh).
Trace Min. Max. Mean Std. Dev. Coeff. of Var. GridStatus
@@
$\triangleright$ Demand Data. To model the data center’s demand, we use Alibaba’s 2018 production trace Alibaba:18, which provides seven days of task start and end times at a 1-second granularity. In our experiments, time steps are fixed to 15 minutes (matching the price data), so we aggregate the demand: we partition the trace into 15-minute buckets and compute a weighted number of tasks for each bucket based on the number of jobs and their active duration (a task that runs for the entire bucket contributes 1, while one that runs for only 5 minutes contributes $\nicefrac{{1}}{{3}}$). We plot this quantity for the full trace in the Appendix, in Fig. 9. Based on this, we compute a scaled demand as $\nicefrac{{\text{weighted no. of tasks}}}{{\text{scale factor}}}$, where a scale factor is chosen based on the size of the energy storage. In most of our experiments, we set this factor to reflect a battery capacity that can meet the daily peak demand for a single (15-minute) time step on $\nicefrac{{2}}{{7}}$ days in the trace. In some experiments, we vary this scale factor to study the effect of storage capacity. Since the trace does not specify task types, we also set a parameter $\texttt{prop\_base}\in[0.5,1]$ to probabilistically classify tasks as base or flexible: with probability prop_base, a task is “base”, and otherwise it is “flexible.” We typically set $\texttt{prop\_base}=0.5$, which yields instances that consist of half base and half flexible demand on average.
@@
$\triangleright$ Algorithms. We implement and test three algorithms, denoted as PAAD, PALD-S, and PALD-C, and solve for the offline optimal solution using GurobiPy gurobi to compute the empirical competitive ratio (ECR). PAAD is the competitive algorithm described in § 4, while PALD-S and PALD-C are variants of the learning framework described in § 6.2 – we typically set a robustness factor of $\rho=5\alpha$ (where $\alpha$ is the standard competitive ratio) to ensure robustness while allowing for some benefit from learning, and set $K=10$. We implement the PALD algorithm using CVXPYLayers CVXPYlayers:19; CVXPY to enable optimization of the downstream task performance via backpropagation. PALD-S directly learns one set of threshold functions to minimize overall ECR – to capture some seasonality in the traces, we train a set of PALD-S thresholds for each month in a given region’s LMP data (i.e., 12 sets of thresholds total), using 100 random training instances for each month. In contrast, PALD-C is a neural network-based approach that uses contextual features to predict a set of threshold functions for each driver – we train it on 100 random training instances using all 12 months in a given region’s LMP data, using the time, month, tracking target (if $\eta>0$), and statistics about the day-ahead LMP forecasts for future time steps (i.e., $\min$, $\max$, average, standard deviation), as features. We defer more details of the PALD implementations to § A.3.
[…]
Figure 3: CDFs of ECR for all algorithms in default CAISO experiments.
@@
Figure 4: Average ECR for all algorithms in varying regions, with a tracking cost, and varying storage size. In (a), top subfigure is coeff. of var. for each region.
$\triangleright$ Grid Regions. To capture the effect of different electricity price distributions and characteristics, 4(a) shows the results of the main experiment with the same parameters in three other regions, namely ERCOT, PJM, and ISONE. Both PALD-S and PALD-C are retrained on a test of training instances generated from each region. We expect to see some variance in performance across these regions, since the global price bounds ($p_{\min}$ and $p_{\max}$) are not homogeneous – interestingly, we find that the performance of algorithms roughly corresponds to the variability of prices in each region, with ERCOT (highest coefficient of variation) showing the most challenging case.
$\triangleright$ Tracking Case. In 4(b), we show algorithm performance in a setting with a tracking cost (i.e., instead of a switching cost, so $\gamma=0$), varying $\eta$ between 0 and $25$. Both PALD-S and PALD-C are retrained with $\eta=10$. We find that the performance of PAAD and PALD-C slightly improves as $\eta$ grows – PALD-S continues to improve on PAAD in the small $\eta$ regime, but for large (out-of-distribution) $\eta>15$, PALD-S’s performance degrades because the single set of thresholds PALD-S learns becomes overly conservative (i.e., waits too long to purchase) when $\eta$ grows.
$\triangleright$ Storage Capacities. Recall that the default setting is where the storage is sized to meet the peak demand on two of the seven days in the Alibaba trace for a single 15-minute time slot – in 4(c), we plot the effect of larger storage sizes that meet up to $4$ hours of the same peak demand on two of the seven days (i.e., a $16\times$ increase in storage size). As the storage size increases and OPT gains flexibility for shifting, the performance of all algorithms slightly degrades, although PALD-C consistently outperforms the other algorithms across all storage cases.
@@
(d) Delivery cost ($c$ & $\varepsilon$)
@@
Figure 5: Average ECR for all algorithms in CAISO with varying parameters as specified.
@@
$\triangleright$ Time Horizons. 5(a) shows the effect of varying the time horizon $T$ between $24$ and $96$ in CAISO. We find that the performance of all algorithms improves slightly with longer horizons, which we attribute to the increased flexibility in using storage over longer time periods. Notably, PALD-C consistently outperforms the other algorithms across all horizon lengths.
@@
$\triangleright$ Switching Costs. 5(b) shows the effect of varying the switching cost coefficient $\gamma$ (with fixed $\delta=5$) between 0 and $25$ in CAISO. As predicted by the theoretical results, the performance of all algorithms degrades slightly as the switching cost coefficient $\gamma$ increases – in this respect, PALD-S and PALD-C are less sensitive to increases in switching costs than PAAD.
@@
$\triangleright$ Varying Demands. We estimate the effect of varying demand distribution on the performance of all algorithms in the CAISO region. Recall that the default setting splits demand evenly between base and flexible demand ($\texttt{prop\_base}=0.5$) – in 5(c), we plot the effect of varying prop_base between $0.5$ and $1.0$ in the CAISO region. We find that the performance of PAAD and PALD-S generally improves as the proportion of base demand increases (deviating further from the training instances), while PALD-C’s performance slightly degrades, although it still outperforms the other algorithms.
@@
$\triangleright$ Delivery Costs. In 5(d), we show the effect of varying the delivery cost parameters $c$ and $\varepsilon$ (see Def. 2.4) in the CAISO region. As $c$ and $\varepsilon$ increase, the delivery cost has a more significant effect on the resulting cost. As predicted by the theoretical results, PAAD’s performance degrades slightly as a function of $c$ and $\varepsilon$, while the performance of PALD-S and PALD-C remains relatively constant.
8 Conclusion

### v2

@@ 6.1. The PALD Framework
To tackle the above challenges, rather than defining a new prediction model and designing an algorithm around it, we propose a new learning framework (Partitioned Accounting & Learned Decisions (PALD)) that directly “learns the best algorithm” from within a general characterized set of algorithms, where each algorithm in the set satisfies a certain desired robustness (i.e., each algorithm guarantees a certain competitive ratio).
@@
Motivation. Under the pseudo-cost minimization framework, a key step in bounding the competitive ratio of PAAD (Thm. 4.3) uses characteristics of the threshold functions ($\phi_{b}$ for base drivers, $\phi_{f}$ and $\psi_{f}$ for flexible drivers) that govern the algorithm’s decisions. In particular, the definition of the pseudo-cost minimization problem allows us to relate PAAD’s cost to a lower bound on OPT via these threshold functions. Then, for a given target robustness $\rho>\alpha$ (where $\alpha$ is the competitive ratio of PAAD for OSDM-S), we can exactly define a feasible set of threshold functions, based on the necessary relations for the competitiveness proof, that guarantee $\rho$-robustness for PALD. A similar characterization of “robust and consistent thresholds” was derived by Benomar et al. (2025) for one-max search to analytically investigate smoothness in algorithms with predictions. In PALD however, our goal is to directly learn the best thresholds within a feasible set that only considers robustness, empirically optimizing performance on historical instances while retaining a bounded competitive ratio. Given learned thresholds $\hat{\phi}_{b},\hat{\phi}_{f}$ and $\hat{\psi}_{f}$, PALD’s pseudocode is exactly the same as that of PAAD (see Algorithm 2) but with $\hat{\phi}_{b},\hat{\phi}_{f}$ and $\hat{\psi}_{f}$ in place of $\phi_{b},\phi_{f}$ and $\psi_{f}$ (analytical thresholds).
@@
$\qquad\Phi_{b}(0,w)+(1-w)(p_{\max}+2\gamma)+p_{\max}(c+\varepsilon)+2\delta-cwp_{\min}\leq\rho\left[\phi_{b}(w)-2\gamma+\varepsilon p_{\max}+\nicefrac{{2\kappa}}{{T}}\right]\big\}.$
@@
$\hskip 87.87442pt\Psi_{f}(0,v)+(1\!-\!v)(p_{\max}(c\!+\!\varepsilon)\!+\!2\delta)\leq\rho\big[\nicefrac{{1}}{{\omega}}\big(\phi_{f}(w)+\psi_{f}(v)-2\kappa\big)+\nicefrac{{2\kappa}}{{T}}\big]\big\}.$
@@
Proof Sketch of Theorem 6.3. (full proof in § G.1) Since PALD is identical to PAAD save for the learned threshold functions, we inherit feasibility from Lemma E.1. First, we lower bound the cost of OPT; as in the proof of Thm. 4.3, we partition the time horizon into active and inactive periods. Since the partitioning relies on the value of the base driver threshold function and PALD’s is learned, we must use the definition of $\mathcal{R}_{b}$ to prove that OPT’s cost during inactive periods can be bounded in the same manner. Next, we upper bound the cost of PALD by “charging” each driver’s decisions to their corresponding threshold functions and accounting for any worst-case purchasing, switching, and delivery costs that arise due to constraints. We show that the same charging argument from Thm. 4.3 still holds with learned thresholds. Finally, we use the definitions of $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ to relate the cost of PALD to that of OPT. Using a similar contradiction argument as in the proof of Thm. 4.3, the last inequalities in the definition of the feasible sets are exactly those needed to show that the ratio $\frac{\texttt{PALD}(\mathcal{I})-p_{\max}\hat{s}}{\texttt{OPT}(\mathcal{I})}$ must be at most $\rho$, completing the proof. $\square$
@@
Remark. PALD is similar to certain existing approaches, but distinct in important ways. For instance, Zeynali et al. (2021) propose learning the best exponent for an exponential threshold function in the online knapsack problem—our approach is distinguished by its expressiveness in allowing any arbitrary threshold function while retaining worst-case guarantees. Our concept of “learning a threshold” is structurally similar to the notion of learning dual functions in online convex optimization (Lobos et al., 2021)PALD is distinguished by the “robustness certificate,” which additionally guarantees a certain (adversarial) competitive ratio. Li et al. (2022) learn an ML model to solve online convex optimization that is calibrated (although not guaranteed) to be robust with respect to a baseline—in contrast, PALD allows us to specify the worst-case level performance we are willing to incur, and to learn an algorithm that performs well subject to that constraint with certainty.
6.2. Learning Methodology
@@
In this section, we describe the methodology we use in our case study (see § 7) to instantiate a “proof-of-concept” of the PALD framework in practice. We parameterize each threshold function as a piecewise-affine function using a “knot” representation at a fixed grid of $K>1$ points $\{0,\nicefrac{{1}}{{K-1}},\nicefrac{{2}}{{K-1}},\ldots,1\}$. We learn a parameter vector $\mathbf{y}\in\mathbb{R}_{+}^{K}$ where each element defines the function’s value at a knot (between knots, the function is interpolated linearly). We denote by $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ the parameter vectors for $\smash{\hat{\phi}_{b},\hat{\phi}_{f},\hat{\psi}_{f}}$, respectively. This makes the robust sets tractable:
@@
The full proof is in § G.3—the key idea is that the necessary robustness certificates for PALD reduce to box constraints, linear inequalities, and affine halfspaces, whose intersection is convex.
@@
Recall that the analytical thresholds $\phi_{b},\phi_{f}$ and $\psi_{f}$ (Defs. 4.1 & 4.2) are continuous and monotone non-increasing functions—by definition of the PAAD algorithm, these lie in the feasible sets $\mathcal{R}_{b}(\alpha)$ and $\mathcal{R}_{f}(\alpha)$, respectively. For $\rho$ increasing away from $\alpha$, the feasible sets $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ become strictly larger, since the robustness conditions become easier to satisfy. Since piecewise-affine functions are known to be dense in continuous functions (Cheney, 1998; Pinkus, 2005), the analytical thresholds can be approximated arbitrarily well using piecewise-affine functions with sufficiently many knots. Thus, for a given robustness $\rho>\alpha$, the feasible sets $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ are non-empty since they at least contain the analytical thresholds. Lemma 6.4 implies that it is efficient to project any learned piecewise-affine function into the feasible sets—this makes robustness tractable to enforce during learning, and unlocks several learning techniques which we discuss below. [^4] In principle, PALD could use substantially different learning techniques than what we consider here—our main contribution is the notion of directly finding “the best algorithm” amongst a set of algorithms satisfying a given worst-case guarantee.
@@
Directly Learning Thresholds. We implement the PALD algorithm by using differentiable convex optimization layers (e.g., CVXPYLayers (Agrawal et al., 2019)) to solve the pseudo-cost minimization problem at each time step. This allows backpropagation through the entire PALD procedure to minimize empirical competitive ratio on historical instances. Given a data set of historical instances $\mathcal{D}=\{\mathcal{I}_{1},\ldots,\mathcal{I}_{N}\}$, we can use this differentiable implementation of PALD to find a single set of parameters $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ that minimizes the empirical competitive ratio on $\mathcal{D}$, using projected gradient descent (PGD) to ensure that, during training, the learned thresholds lie in the robust sets $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ for a target $\rho>\alpha$. We refer to this as PALD-S (“S” for simple) in our case study.
@@
End-to-End Contextual Learning. While directly learning a single set of thresholds improves average-case performance, it cannot leverage instance-specific context that is extremely useful in practice. For instance, our case study in § 7 uses locational marginal price (LMP) data for different grids; LMPs are known to correlate with factors such as the time of day, the weather, and the season (Qu et al., 2024). Using a learning method that is aware of such context can enable better performance in practice by learning thresholds that are tailored to likely operating conditions.
@@
To realize this context-aware approach, we employ a neural network $F_{\Theta}(\mathbf{x})\to\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ (with parameters $\Theta$) that takes as input a context vector $\mathbf{x}$ about a given driver and outputs predicted vectors $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ that define the threshold functions. The final layer of this neural network is a differentiable projection layer that projects the output parameters $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi}$ into the feasible sets $\mathcal{R}_{b}(\rho)$ and $\mathcal{R}_{f}(\rho)$ to ensure target robustness $\rho>\alpha$. We construct an unsupervised learning pipeline that simulates the differentiable PALD algorithm on a historical data set $\mathcal{D}$. For each instance, each new driver in PALD makes a forward pass through the neural network $F_{\Theta}$ to obtain its context-aware thresholds $\smash{\hat{\phi}_{b},\hat{\phi}_{f}}$, and $\smash{\hat{\psi}_{f}}$. After observing the resulting performance, we backpropagate through the entire pipeline to learn a better mapping between context and thresholds. We refer to this technique as PALD-C in our case study, and give a diagram of it in Appendix A.3.
@@
7. Case Study: A Grid-integrated Data Center with Local Energy Storage
@@
Table 1. Statistics for LMP traces. All traces span the time period 01/01/2024-12/31/2024 at a 15-minute granularity, and LMP data is reported in \$ (USD) per megawatt-hour (MWh).
Trace Min. Max. Mean Std. Dev. Coeff. of Var. (Grid Status, 2025)
@@
$\triangleright$ Demand Data. To model the data center’s demand, we use Alibaba’s 2018 production trace (Alibaba, 2018), which provides seven days of task start and end times at a 1-second granularity. In our experiments, time steps are fixed to 15 minutes (matching the price data), so we aggregate the demand: we partition the trace into 15-minute buckets and compute a weighted number of tasks for each bucket based on the number of jobs and their active duration (a task that runs for the entire bucket contributes 1, while one that runs for only 5 minutes contributes $\nicefrac{{1}}{{3}}$). We plot this quantity for the full trace in the Appendix, in Fig. 9. Based on this, we compute a scaled demand as $\nicefrac{{\text{weighted no. of tasks}}}{{\text{scale factor}}}$, where a scale factor is chosen based on the size of the energy storage. In most of our experiments, we set this factor to reflect a battery capacity that can meet the daily peak demand for a single (15-minute) time step on $\nicefrac{{2}}{{7}}$ days in the trace. In some experiments, we vary this scale factor to study the effect of storage capacity. Since the trace does not specify task types, we also set a parameter $\texttt{prop\_base}\in[0.5,1]$ to probabilistically classify tasks as base or flexible: with probability prop_base, a task is “base”, and otherwise it is “flexible.” We typically set $\texttt{prop\_base}=0.5$, which yields instances that consist of half base and half flexible demand on average.
@@
$\triangleright$ Algorithms. We implement and test three algorithms, denoted as PAAD, PALD-S, and PALD-C. We solve for the offline optimal solution using GurobiPy (Gurobi Optimization, LLC, 2024) to compute the empirical competitive ratio (ECR). PAAD is the competitive algorithm described in § 4, while PALD-S and PALD-C are instantiations of the learning framework described in § 6—we typically set a robustness factor of $\rho=5\alpha$ (where $\alpha$ is the standard competitive ratio) to ensure robustness while allowing for some benefit from learning, and set $K=10$. We implement the PALD algorithm using CVXPYLayers (Agrawal et al., 2019; Diamond and Boyd, 2016) to optimize downstream task performance via backpropagation. PALD-S directly learns one set of threshold functions to minimize overall ECR—to capture seasonality in the traces, we train a set of PALD-S thresholds for each month in a given region’s LMP data (i.e., 12 sets of thresholds total), using 100 random training instances for each month. In contrast, PALD-C is a neural network-based approach that uses context features to predict threshold functions for each driver—we train it on 100 random training instances using all 12 months in a given region’s LMP data, using the time, month, tracking target (if $\eta>0$), and statistics about day-ahead LMP forecasts (i.e., $\min$, $\max$, avg., std. dev.), as features. We defer more implementation details of PALD-S & PALD-C to § A.3.
Figure 3. CDFs of ECR for all algorithms in default CAISO experiments.
@@
Figure 4. Average ECR for all algorithms in varying regions, with a tracking cost, and varying storage size. In (a), top subfigure is coeff. of var. for each region.
[…]
$\triangleright$ Grid Regions. To capture the effect of different electricity price distributions and characteristics, Fig. 4(a) shows the results of the main experiment with the same parameters in three other regions, namely ERCOT, PJM, and ISONE. Both PALD-S and PALD-C are retrained on a test of training instances generated from each region. We expect to see some variance in performance across these regions, since the global price bounds ($p_{\min}$ and $p_{\max}$) are not homogeneous—interestingly, we find that the performance of algorithms roughly corresponds to the variability of prices in each region, with ERCOT (highest coefficient of variation) showing the most challenging case.
$\triangleright$ Tracking Case. In Fig. 4(b), we show algorithm performance in a setting with a tracking cost (i.e., instead of a switching cost, so $\gamma=0$), varying $\eta$ between 0 and $25$. Both PALD-S and PALD-C are retrained with $\eta=10$. We find that the performance of PAAD and PALD-C slightly improves as $\eta$ grows—PALD-S continues to improve on PAAD in the small $\eta$ regime, but for large (out-of-distribution) $\eta>15$, PALD-S’s performance degrades because the single set of thresholds PALD-S learns becomes overly conservative (i.e., waits too long to purchase) when $\eta$ grows.
$\triangleright$ Storage Capacities. Recall that the default setting is where the storage is sized to meet the peak demand on two of the seven days in the Alibaba trace for a single 15-minute time slot—in Fig. 4(c), we plot the effect of larger storage sizes that meet up to $4$ hours of the same peak demand on two of the seven days (i.e., a $16\times$ increase in storage size). As the storage size increases and OPT gains flexibility for shifting, the performance of all algorithms slightly degrades, although PALD-C consistently outperforms the other algorithms across all storage cases.
@@
(d) Delivery cost ($c$ and $\varepsilon$)
@@
Figure 5. Average ECR for all algorithms in CAISO with varying parameters as specified.
@@
$\triangleright$ Time Horizons. Fig. 5(a) shows the effect of varying the time horizon $T$ between $24$ and $96$ in CAISO. We find that the performance of all algorithms improves slightly with longer horizons, which we attribute to the increased flexibility in using storage over longer time periods. Notably, PALD-C consistently outperforms the other algorithms across all horizon lengths.
@@
$\triangleright$ Switching Costs. Fig. 5(b) shows the effect of varying the switching cost coefficient $\gamma$ (with fixed $\delta=5$) between 0 and $25$ in CAISO. As predicted by the theoretical results, the performance of all algorithms degrades slightly as the switching cost coefficient $\gamma$ increases—in this respect, PALD-S and PALD-C are less sensitive to increases in switching costs than PAAD.
@@
$\triangleright$ Varying Demands. We estimate the effect of varying demand distribution on the performance of all algorithms in the CAISO region. Recall that the default setting splits demand evenly between base and flexible demand ($\texttt{prop\_base}=0.5$)—in Fig. 5(c), we plot the effect of varying prop_base between $0.5$ and $1.0$ in the CAISO region. We find that the performance of PAAD and PALD-S generally improves as the proportion of base demand increases (deviating further from the training instances), while PALD-C’s performance slightly degrades, although it still outperforms the other algorithms.
@@
$\triangleright$ Delivery Costs. In Fig. 5(d), we show the effect of varying the delivery cost parameters $c$ and $\varepsilon$ (see Def. 2.4) in the CAISO region. As $c$ and $\varepsilon$ increase, the delivery cost has a more significant effect on the resulting cost. As predicted by the theoretical results, PAAD’s performance degrades slightly as a function of $c$ and $\varepsilon$, while the performance of PALD-S and PALD-C remains relatively constant.
8. Conclusion